A gaseous mixture of He and 02 is found to have a density of 0.518 g/L at 25°C and 720 torr. What is the % by mass of helium in the mixture

Grade:

1 Answers

Arun
5 years ago
First, you can use the ideal gas law to calculate the average molar mass of the gas mixture: 

PV = n RT 
Since n = mass / molar mass 
PV = (mass/molar mass) RT 

Since density = mass/V, you can rearrange the equation to give: 

molar mass = (density)RT / P 

Using the values in the problem (and converting everything to the correct units) 

molar mass = (0.518 g/L) (0.0821 Latm/molK) (298K) / 0.9487 atm 
molar mass = 13.35 g/mol 

Now, let x = mole fraction of He and 1-x = mole fraction of O2, Then: 

13.35 g/mol = 4(x) + 32(1-x) 

Once you solve for x, just multiply it by 100 to get the percent of helium in the gas mixture.